A money manager is looking at various investment options. Her clients are looking to invest some excess cash in a low-risk investment and do not anticipate needing this cash for a minimum of two years. Currently, the manager can invest her clients' money for one year at an effective annual rate of 3.889% with the intention of reinvesting the money for an additional year at whatever prevailing one-year spot interest rates are one year from today. Alternatively, she can make a two-year investment and lock in an effective annual rate of 4.163%.

What is the implied one-year forward rate for year two at which the money manager, acting on her clients' behalf, should be indifferent between the two options? (Given the one- and two-year spot rates provided, what is the implied one-year spot rate, one year from now?)
Answer #1

implied one-year spot rate, one year from now=(1+rate for 2 years)^2/(1+rate for 1 year)-1=1.04163^2/1.03889-1=4.4377227%